you have 2 liters of a 10% solution. how much of a 25% solution would you have to add to get a 16% solution?

Answer by TimothyLamb(4379)

10*2 + 25x = 16(2 + x)
20 + 25x = 32 + 16x
25x - 16x = 32 - 20
9x = 12
x = 12/9
x = 1.33333333333 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php